OpenStudy (anonymous):

Solve : 3^4x-7= 4^2x+3

OpenStudy (anonymous):

this a repeat?

myininaya (myininaya):

Yep it is...let me find it

OpenStudy (anonymous):

start with \[(4x-7)\ln(3)=(2x+3)\ln(4)\] then algebra

OpenStudy (anonymous):

\[4\ln(3)x-7\ln(3)=2\ln(4)x-3\ln(4)\] \[4\ln(3)x-2\ln(4)x=7\ln(3)-3\ln(4)\] \[(4\ln(x)-2\ln(4))x = 7\ln(3)-3\ln(4)\] \[x=\frac{7\ln(3)-3\ln(4)}{4\ln(x)-2\ln(4)\]

OpenStudy (anonymous):

\[x=\frac{7\ln(3)-3\ln(4)}{4\ln(x)-2\ln(4)}\]